OpenStudy (anonymous):

Verify the identity. 4 csc 2x = 2 csc^2(x) tan x

OpenStudy (anonymous):

help me!!!!:(

OpenStudy (anonymous):

hi there

OpenStudy (anonymous):

sure simple

OpenStudy (anonymous):

@mathsmind hi!!

OpenStudy (anonymous):

ok now step one

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

sin2(x)=2sin(x)cos(x)

OpenStudy (anonymous):

sorry be patient my net is slow

OpenStudy (anonymous):

csc(x) = 1/ sin(x)

OpenStudy (anonymous):

so 4csc(2x) = 4/sin(2x)

OpenStudy (anonymous):

you agree so far

OpenStudy (anonymous):

4/sin(2x)=2/sin(x)cos(x)

OpenStudy (anonymous):

now am going to use the equation editor to make it clearer for u ok

OpenStudy (anonymous):

we are going to multiply both sides by sin(x)/sin(x)

OpenStudy (anonymous):

okay!!

OpenStudy (anonymous):

\[=\frac{ 2 }{ \sin(x)\cos(x) }\frac{ \sin(x) }{ \sin(x) }\]

OpenStudy (anonymous):

\[==\frac{ 2 }{ \sin^2(x) }\frac{ \sin(x) }{ \cos(x) }\]

OpenStudy (anonymous):

tan(x) = sin(x)/cos(x), and 2/sin^(x) = 2csc^2(x)

OpenStudy (anonymous):

\[=2 \tan(x)\csc^2(x)\]

OpenStudy (anonymous):

this proves your claim

OpenStudy (anonymous):

i hope this helped

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

thankyou!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

you are welcome!

OpenStudy (anonymous):

you need to understand and mesmerism all tig identities like ur name if you don't want to suffer on this topic

OpenStudy (anonymous):

memorize*